3(2/3x-4)=9
Please try to explain what you're doing. I dont understand how to get the answer. Thanks

Answer by algebrahouse.com(1659)

"3(2/3x-4)=9
Please try to explain what you're doing. I dont understand how to get the answer. Thanks"

3(2/3x - 4) = 9
3(2/3x) + 3(-4) = 9 {used distributive property with 3, multiplying by each term in parentheses}
2x - 12 = 9 {multiplied through}
2x = 21 {added 12 to both sides}
x = 21/2 {divided both sides by 2}
x = 10 1/2 {changed to mixed number}
www.algebrahouse.com